Unlevered firm Levered firm
EBIT 10000 10000
Interest 0 3200
Taxable income 10000 6800
Tax (tax rate: 34%) 3400 2312
Net income 6600 4488
CFFA 6600 7688

The firm is originally 100% financed by equity (Unlevered firm). Assuming that cost of debt =8%; unlevered cost of capital =10%; tax rate= 34%; systematic risk of the asset is 2. Assuming that the firm issues $ 40,000 to buy back some shares, and the debts are traded at par value.

a) What is the annual of the tax shield?

b) Calculate from the unlevered firm: 1) Value of the unlevered firm , 2) Equity for the unlevered firm

c) Suppose that the firm changes its capital structure so that the debt-to-equity ratio is 2.0, then recalculate the following:1) cost of equity, 2) cost of capital, 3) systematic risk of the equity
Answer #1

a) annual of tax shield is under : 6600 @ 34% is =2244 (unlevraged firm)

annual of tax shied is under : 4488 @ 34%is = 1525.92(levraged firm)

b) 1.value of unlevered firm is as under : 6600+40000=$ 46600

2. equity is : 6600+40000= $ 46600

(c) cost of equity : 46600/3 i.e = 15533.33

15533.33/46600*100=33.33-12=21.33%

cost of capital =21.33+12=33.33%

systamatic risk of equity is = 2/2*100= 1

Answer